Find the domain and range of the following rational function. Use qny notation. 1) f(x) = 2 / x+1 2) f(x) = 3x / x+3 3) f(x) = 3-x / x-7 4) f(x) = 2+x / x 5) f(x) = (x+1) / x^2-1
1
Expert's answer
2021-10-26T17:29:49-0400

1) f(x) = "\frac{2}{x+1}"


Here x+1 "\not =" 0 // otherwise function will not be defined

So x "\not =" - 1


Domain ("-\infty,-1)" "\bigcup" "(-1 , \infty)"


Range "(-\infty ,0) \bigcup (0,\infty)"



2) f(x) = "\frac{3x}{x+3}"


Here x+3 "\not =" 0 // otherwise function will not be defined

So x "\not =" -3


Domain "(-\infty ,-3) \bigcup (-3 ,\infty)"


f(x) = "\frac{3x +9 -9}{x+3}" = "\frac{3(x+3) - 9}{x+3}" = 3 - "\frac{9}{x+3}" ----------------------- (1)


From equation (1) we can understand the value will never reach 3.


So Range "(-\infty ,3) \bigcup (3, \infty)"


3) f(x) = "\frac{3-x}{x-7}"


Here x-7 "\not =" 0 // otherwise function will not be defined


So x "\not =" 7


Domain "(-\infty ,7) \bigcup (7,\infty)"


f(x) = "\frac{-(x-3)}{x-7}" = "\frac{-(x-3-4+4)}{(x-7)}" ="\frac{-(x-7+4)}{(x-7)}" = -1 +"\frac{-4}{(x-7)}"


Range "(-\infty ,-1) \bigcup (-1,\infty)"



4) f(x) = "\frac{2+x}{x}"


Here x "\not =" 0


Domain "(-\infty ,0) \bigcup (0,\infty)"


f(x) = "\frac{2+x}{x}" = "\frac{2}{x}" + 1


Range "(-\infty,1) \bigcup (1,\infty)"


5) f(x) = "\frac{x+1}{x^2 -1}"


Here x2 - 1 "\not =" 0

So x "\not = 1" and x "\not = -1"


Domain "(-\infty,-1) \bigcup (-1,1) \bigcup (1,\infty)"


Range "(-\infty ,0) \bigcup (0,\infty)"